Q1. The length of a rectangular swimming pool is 2x-1 meters, and the width is x + 2 meters. Write a polynomial A(x) that represents the area. Find A(5).
Q2.Find a trinomal A(x) that represents the area of a parallelogram whose base is 3x + 2 meters and whose height is 2x + 3 meters. Find A(3).
Q3.The perimeter of a rectangular backyard is 6x +6 yards. If the width is x yards, find a binomial that represents the length.
Q4.P dollars is inveted at annual interest rate r for 1 year. If the interest is compounded semiannually, then the polynomial P(1 + r/2)to the 2nd power represents the value of the investment after 1 year. Rewrite this expression without parentheses. Evaluate the polynomial if P=$200 and r=10%.

Answer by stanbon(75887)

Q1. The length of a rectangular swimming pool is 2x-1 meters, and the width is x + 2 meters. Write a polynomial A(x) that represents the area. Find A(5).
--------
A(x) = (2x-1)(x+2)
A(5) = (10-1)(5+2) = 9*7 = 63
---------------------------------

Q2.Find a trinomal A(x) that represents the area of a parallelogram whose base is 3x + 2 meters and whose height is 2x + 3 meters. Find A(3).
A(x) = (3x+2)(2x+3)
A(3) = (9+2)(6+3) = 11*9 = 99
---------------------------------

Q3.The perimeter of a rectangular backyard is 6x +6 yards. If the width is x yards, find a binomial that represents the length.
Perimeter = 2(L+w)
6x+6 = 2(L + x)
3x+3 = L + x
Length = 2x+3
---------------------------------

Q4. P dollars is invested at annual interest rate r for 1 year. If the interest is compounded semiannually, then the polynomial P(1 + (r/2))to the 2nd power represents the value of the investment after 1 year. Rewrite this expression without parentheses. Evaluate the polynomial if P=$200 and r=10%.
---
P(1+(r/2))^2 = P(1 + r + (r/2)^2)
---
200(1 + 0.10 + (0.05)^2)
= 200(1 + 0.10 + 0.0025)
= 200(1.1025)
= 220.50
===========================
Cheers,
Stan H.